OpenStudy (anonymous):

The bones of an average adult person weigh approx. 10.9g and contain approx. 48% Ca3(PO4)2. Calculate the mass of phosphorus present in the skeleton of the average adult.

OpenStudy (matt101):

(You probably mean 10.9 kg = 10900 g for the mass of the bones in an average adult) 48% of the bone is composed of calcium phosphate, so of the entire bone mass, 10900*0.48 = 5232 g is calcium phosphate. To find the mass of phosphorus in the skeleton, first find the mass of phosphorus in one "molecule" of Ca3(PO4)2. The molar mass calcium phosphate is 310 g, 120 g from the calcium and 190 g from the phosphate. This means 190/310 = 61.3% of the mass of calcium phosphate comes from just the phosphate. If you have 5232 g of calcium phosphate, 61.3%, or 3207 g comes from phosphate and this is how much is present in the skeleton. If you have any questions please let me know!

OpenStudy (anonymous):

thanks a lot! how di you get 3207g from 61.3%?

OpenStudy (matt101):

5232 * 0.613 = 3207